I need your help onthe following problem: "A rectangle whose perimeter is 250 is 8 feet shorter than it is wide.Find the dimensions and area." Can you please help me.
Answer by ankor@dixie-net.com(12701)

A rectangle whose perimeter is 250 is 8 feet shorter than it is wide.Find the dimensions and area." Can you please help me.
:
It says: 2L + 2W = 250
:
It also says: L = W - 8
:
Sustitute (W-8) for L in the perimeter equation:
:
2(W-8) + 2W = 250
2W - 16 + 2W = 250
4W = 250 + 16
4W = 266
W = 266/4
W = 66.5 ft is the width
:
L = 66.5 - 8 = 58.5 ft is the length
:
Check our solutions in the perimeter equation:
2(58.5) + 2(66.5) =
117 + 133 = 250
:
Find the area: 58.5 * 66.5 = 3890.25 sq ft
:
Make sense to you??